# In Situ Training of Implicit Neural Compressors for Scientific Simulations via Sketch-Based Regularization

### v2

@@ 2.1 Sketching
Sketches may be deterministic or randomized. Often, in the latter case, we use random matrix ensembles based on specific distributions such as the Gaussian. Random subsampling can also be cast in this form by choosing a random subset of the $n$-dimensional standard basis. Gaussian ensembles generally perform well and have a plethora of theoretical results, but require slow, dense matrix operations. On the other hand, subsampling is extremely fast but can perform quite poorly in specific scenarios. In the worst-case, subsampling may select points that are the least helpful for the desired task, whereas dense sketches always aggregate information from all points. However, subsampling does preserve part of the mesh structure, as it is coarsening instead of aggregating, which can be beneficial in certain settings. Deterministic sketches with this preservation property have been studied in [dunton2021deterministic], with the idea that sketching operators informed by the data itself can outperform, both in speed and approximation quality, randomized sketches which are usually data oblivious.
@@
We elect to use the fast Johnson-Lindenstrauss transform (FJLT) [ailon2009fast] due to its superior speed and empirical performance compared to other possible options. However, we will also investigate the performance of subsampling in Section˜3 — and it is also used as a subroutine for the FJLT sketch — so we present this in Algorithm˜2. The FJLT is a Johnson-Lindenstrauss transform (JLT): a transformation that projects a finite set of points into many fewer dimensions while approximately preserving pairwise relative distances. The JL lemma (Theorem˜1) proves the existence of such a mapping, and any such mapping that satisfies this lemma may be referred to as a JL transform.
Theorem 1 (Johnson-Lindenstrauss Transformations [johnson1984extensions]).
@@
Furthermore, the construction of the distribution $\mathcal{D}$ is explicit, with the JLT first implemented via an orthonormalized Gaussian random matrix, for which Eq.˜5 holds with $k=\mathcal{O}(\epsilon^{-2}\log(m))$ [johnson1984extensions]. The practical downside is the slow, dense matrix-vector multiplications. The FJLT, on the other hand, uses a significantly different approach to gain a similarly significant speedup. Algorithm˜3 describes this process, which uses the orthogonal discrete cosine transform (DCT) to effectively mix rows [avron2010blendenpik] [^3] Specifically, a DCT of type II or III.. Algorithm˜3 satisfies Eq.˜5 with $k=\mathcal{O}(\epsilon^{-2}\log(m(\delta-n^{-\log^{3}(n)})^{-1})\log^{4}(n))$ (see, e.g., [huang2021spectral]) and, due to the DCT, only incurs an $\mathcal{O}(n\log(n))$ time cost since the DCT can be applied using fast Fourier transform techniques. As noted earlier, the random seed for the Rademacher random variable and subsampling are saved so that the same FJLT transform can be reused later.
@@
$\Omega\leftarrow\operatorname{random\_choice}(\{1,\ldots,n\},k)$ ;
@@
$\bm{d}\leftarrow\operatorname{rademacher}(n)\in\{\pm 1\}^{n}$ ;
@@
 $\widehat{\bm{U}}_{:,c}\leftarrow\sqrt{n/k}\operatorname{DCT}(\bm{U}_{:,c}\odot\bm{d})$ ;
@@
return $\operatorname{Subsampling-Sketch}(\widehat{\bm{U}},k)$ ;
@@ 2.2 JL Regularization for In Situ Learning
We refer to [Lee_manifoldBook2ndEd] for an extensive treatment of manifolds, but it is sufficient for our purposes to understand that a Riemannian manifold can be viewed as a subset of Euclidean space and behaves locally like a vector space. Informally, we are considering (possibly) lower-dimensional geometric structures embedded in some ambient subset of $\mathbb{R}^{N}$.
Theorem 2 (Manifold JL Transform [baraniuk2009random]).
@@
The volume, condition number, and geodesic covering regularity (the latter two of which are defined in [baraniuk2009random]) together define the topological regularity of the manifold. Note that this result, as originally stated, includes a factor of $\sqrt{k/N}$ as the sketch is not assumed to be normalized. Theorem˜2 presents a very similar result to Theorem˜1, but where the size of the sketch depends on geometric properties of the manifold, most importantly the intrinsic dimension, and not on the collection of points being sketched. This means that for low-dimensional manifolds of a potentially high-dimensional ambient space, one can construct a small JL transform for an arbitrarily large number of points. The size $k$ depends only logarithmically on the ambient dimension $N$, and in fact, this dependence can be removed completely as shown in subsequent works [ArminMike2015]. This ability to construct JL transformations for manifolds is key to our main result, which we present next.
@@
Define $\mathcal{L}$ as the squared $\ell_{2}$ (i.e., unnormalized MSE) loss. For a fixed mesh $\bm{X}$ and time $t\in[T]$, with parameters $\bm{\theta}\in\mathbb{R}^{N}$, suppose the map $\bm{\Phi}:\mathbb{R}^{N}\to\mathcal{M}\subset\mathbb{R}^{n\times c}$, defined via the neural network as $\bm{\Phi}(\bm{\theta})=\bm{\Phi}(\bm{X},t;\bm{\theta})$, maps to the Riemannian manifold $\mathcal{M}$. Further, suppose the sketch $\bm{S}_{t}:\mathbb{R}^{n\times c}\to\mathbb{R}^{k\times c}$ is a $(\epsilon,\delta)$-JL transform on the shifted manifold $\mathcal{M}-\{\bm{U}_{t}\}$, then, $\forall\bm{\theta}\in\mathbb{R}^{N}$, the following holds with probability at least $1-\delta$:
@@
The proof of the second inequality is motivated by the informal proof of Theorem 2.14 in [woodruff2014sketching]. Also, note that we construct a JL transform for the shifted manifold $\mathcal{M}-\{\bm{U}_{t}\}$, which is just a convenience to simplify the proof, but does not introduce any complexity to the sketch itself. The shifted manifold will have the same geometric properties as $\mathcal{M}$ that impact the construction of the sketch.
@@ 3 Experiments
We conduct our experiments on the datasets listed in Table˜1, which highlight several potentially challenging or important attributes. The Ignition dataset describes a fully resolved gas wave front on a 2D uniform grid that transitions from a transport phase to a steady state jet. The Neuron dataset describes a 3D diffusion process in an unstructured, non-Cartesian branching neuron tree [Angran3D]. The Channel dataset, from the Johns Hopkins Turbulence Database [channel_flow_1, channel_flow_2], describes 3D turbulent flow on a non-uniform grid. It has been trimmed in space to a $64^{3}$ volume and in time to the first 500 snapshots. We define the notion of a “sample factor” as the relative size of the sketch to the mesh size for a given dataset, i.e., $100\cdot k/n$.
@@
Equation˜6 provides the frame-wise relative loss function we use for training, Eq.˜7 defines the full dataset relative Frobenius error, and Eq.˜8 defines the frame-wise peak signal to noise ratio (PSNR). The latter two of these three serve as our primary performance metrics.
@@
$\text{PSNR}(\bm{U}_{t},\bm{\widetilde{U}}_{t})=\frac{1}{C}\sum_{c=1}^{C}20\cdot\log_{10}\left(\frac{\max_{i}(\bm{\widetilde{U}}_{t})_{i,c}}{\|(\bm{U}_{t})_{:,c}-(\bm{\widetilde{U}}_{t})_{:,c}\|_{2}}\right).$ (8)
@@
Our work is implemented in PyTorch [pytorch] and available as an open-source GitHub repository [paper-repo]. All training is conducted on a single GPU (A100 or L40S) using the RAdam optimizer, a learning rate of $10^{-4}$, and in single precision (i.e., 32-bit floating point) [single_precision] for both the parameters and data. Given that our goal is to compress the given data, we do not seek to generalize to unseen data; thus, overfitting is not an issue. Our method is fully in situ, so we perform no pre-processing of the data. For all in situ results, the size of the full buffer is always $T_{f}=1$ (only one full snapshot available at a time), and the size of the sketch buffer is $T_{s}=T-1$ (one less than the full time horizon). We do not require sketching the final snapshot because the training loop ends after the last full snapshot has been encoded. The batch size of the sketch buffer and the number of epochs per full snapshot vary between datasets, but are in the range 25–45 and 300–500, respectively.
@@ 3.1 Model Architecture
As alluded to previously, we elect to use a hypernetwork approach for constructing our neural compressor. Practically, this results in two networks working together to map space-time coordinates to a reconstruction of the PDE solution field. Figure˜4 displays the generic INR architecture we employ for both the hypernetwork and the target network. The former maps from just the time coordinate $t$ to the parameters $\bm{\theta}(t)$. The latter ingests $\bm{\theta}(t)$ to define its layers and maps from the space-time coordinates $(\bm{x},t)$ to the reconstruction $\tilde{\bm{u}}(\bm{x},t)$.
@@
Within the sine-layer blocks, we use sinusoidal activation functions with the initialization developed in [sitzmann2020implicit]. We also employ skip connections similar to those of [pan2023neural, sales2024implicit]. Thus, our networks are defined by two hyperparameters: the depth, i.e., the number of blocks, and the width of the internal linear layers. Because we are targeting the in situ setting, we cannot effectively normalize the data to ensure it is in some specific range, so no activation is used at the final linear layer. This ensures that the target network can produce any signal value. The lack of activation is also useful for the hypernetwork, as we construct the initial output of the hypernetwork to match the initialization of the target INR. To accomplish this, we use the strategy of [han2023kd], wherein the bias of the final linear layer in the hypernetwork is set to the exact values given by the initialization of the target INR, and the weights of the last linear layer are scaled by a relatively small factor.
@@ 3.2 Results
Table 2: Performance metrics on test datasets Ignition, Neuron, and Channel at 5%, 1%, and 2% sketch sample factors (where relevant), respectively. Metric statistics are reported as mean plus-or-minus standard deviation from five identical trials. The numbers in the parentheses of the first row of each table are compression ratios.
@@
The results also show that our in situ methods, which combine training of the full snapshot with regularization from previous sketched snapshots, outperform the Offline-FJLT strategy (which just uses the sketched snapshots). The Offline-FJLT is conceptually simple to implement, since it does the sketching in situ and then trains offline. However, since it cannot train on the unsketched data, it incurs the $\frac{1+\epsilon}{1-\epsilon}$ error. Our in situ methods use sketched data but do not seem to incur this extra error, since the sketching is only used to prevent forgetting. This is also unsurprising, since for human memory, the task of learning something new is usually more difficult than the task of not forgetting something already learned.
In Table˜3, we explore how well the theory from Section˜2.2 aligns with our empirical results. Although unreported in Table˜2, we use Offline-FJLT and Offline-Baseline results for Eq.˜6 to compute the $\frac{1+\epsilon}{1-\epsilon}$ ratio from Theorem˜3, which we subsequently use to solve for $\epsilon$. We also compute principal component analysis (lPCA [fukunaga1971algorithm]) based estimates of the local manifold dimension for the INC networks used with each of the target datasets. This local dimensionality estimate yields a rough idea of the overall manifold dimension. We perform this analysis using the scikit-dimension package [scikit-dimension] by taking perturbations of size $10^{-5}$ around a nominal point of network parameters. This nominal state is obtained within the in situ setting after training on the first snapshot, i.e., the first time sketching is used.
@@
Figure 6: Reconstruction comparison of the Channel dataset on all three channel flow velocities ($x,y,z$ components in top to bottom order) at snapshot $t=250$ for a constant $z$ coordinate in the center of the volume.
Per-snapshot relative error (Eq.˜6) curves for training and testing on the Ignition dataset are given in Fig.˜7. The training loss is split according to Eq.˜4 into the two components $\mathcal{L}_{\text{full}}$ and $\mathcal{L}_{\text{sketch}}$ corresponding to the two buffers. Thus, the “Full Loss” can be viewed as the instantaneous loss at the current snapshot, while the “Sketch Loss” is the averaged loss over previous sketched snapshots. The “Test Loss” is simply the instantaneous loss for a fixed snapshot after the network has been trained. For early snapshots, before approximately $t=50$, one can observe that the testing error is significantly better than the training loss, indicating the sketched snapshots have also actively been used for learning instead of only preventing forgetting.
Figure 7: Training and testing error on the Ignition data for the InSitu-FJLT case. Training error is taken at the end of the optimizer steps for a given snapshot. Data has been smoothed via convolution with a Gaussian.
@@
One may ask whether the inclusion of the hypernetwork in our overall INC architecture, shown in Fig.˜4, is necessary? Indeed, sketching as a regularization is just as applicable to a time-dependent INR, which we have found to be just as expressive in the offline setting (see Fig.˜2). To probe this question, Table˜4 presents comparison results on the Ignition dataset in the in situ regime with and without a hypernetwork. In the latter case, the hypernetwork branch seen in Fig.˜4(b) is simply eliminated. We include results for both our subsample and FJLT sketching options. Eliminating the hypernetwork results in fewer parameters, so the “Hypernet + INR” model is appropriately downsized to ensure approximately the same compression rate. We observe that the addition of a hypernetwork significantly increases performance regardless of the sketch.

### v3

@@ 2.1 Sketching
Sketches may be deterministic or randomized. Often, in the latter case, we use random matrix ensembles based on specific distributions such as the Gaussian. Random subsampling can also be cast in this form by choosing a random subset of the $n$-dimensional standard basis. Gaussian ensembles generally perform well and have a plethora of theoretical results, but require slow, dense matrix operations. On the other hand, subsampling is extremely fast but can perform quite poorly in specific scenarios. In the worst-case, subsampling may select points that are the least helpful for the desired task, whereas dense sketches always aggregate information from all points. However, subsampling does preserve part of the mesh structure, as it is coarsening instead of aggregating, which can be beneficial in certain settings. Deterministic sketches with this preservation property have been studied in [14], with the idea that sketching operators informed by the data itself can outperform, both in speed and approximation quality, randomized sketches which are usually data oblivious.
@@
We elect to use the fast Johnson-Lindenstrauss transform (FJLT) [1] due to its superior speed and empirical performance compared to other possible options. However, we will also investigate the performance of subsampling in Section˜3 — and it is also used as a subroutine for the FJLT sketch — so we present this in Algorithm˜2. The FJLT is a Johnson-Lindenstrauss transform (JLT): a transformation that projects a finite set of points into many fewer dimensions while approximately preserving pairwise relative distances. The JL result Theorem˜1 proves the existence of such a mapping, and any such mapping that satisfies this lemma may be referred to as a JL transform.
Theorem 1 (Johnson-Lindenstrauss Transformations [27]).
@@
Furthermore, the construction of the distribution $\mathcal{D}$ is explicit, with the JLT first implemented via an orthonormalized Gaussian random matrix, for which Eq.˜5 holds with $k=\mathcal{O}(\epsilon^{-2}\log(m))$ [27]. The practical downside is the slow, dense matrix-vector multiplications. The FJLT, on the other hand, uses a significantly different approach to gain a similarly significant speedup. Algorithm˜3 describes this process, which uses the orthogonal discrete cosine transform (DCT) to effectively mix rows [3] [^4] Specifically, a DCT of type II or III.. Algorithm˜3 satisfies Eq.˜5 with $k=\mathcal{O}(\epsilon^{-2}\log(m(\delta-n^{-\log^{3}(n)})^{-1})\log^{4}(n))$ (see, e.g., [25]) and, due to the DCT, only incurs an $\mathcal{O}(n\log(n))$ time cost since the DCT can be applied using fast Fourier transform techniques. As noted earlier, the random seed for the Rademacher random variable and subsampling are saved so that the same FJLT transform can be reused later.
@@
$\Omega\leftarrow\operatorname{random\_choice}(\{1,\ldots,n\},k)$
@@
$\bm{d}\leftarrow\operatorname{rademacher}(n)\in\{\pm 1\}^{n}$
@@
 $\widehat{\bm{U}}_{:,c}\leftarrow\sqrt{n/k}\operatorname{DCT}(\bm{U}_{:,c}\odot\bm{d})$
@@
return $\operatorname{Subsampling-Sketch}(\widehat{\bm{U}},k)$
@@ 2.2 JL Regularization for In Situ Learning
We refer to [29] for an extensive treatment of manifolds, but it is sufficient for our purposes to understand that a Riemannian manifold can be viewed as a subset of Euclidean space and behaves locally like a vector space. Informally, we are considering (possibly) lower-dimensional geometric structures embedded in some ambient subset of $\mathbb{R}^{N}$.
Theorem 2 (Manifold JL Transform [5]).
@@
The volume, condition number, and geodesic covering regularity (the latter two of which are defined in [5]) together define the topological regularity of the manifold. Note that this result, as originally stated, includes a factor of $\sqrt{k/N}$ as the sketch is not assumed to be normalized. Theorem˜2 presents a very similar result to Theorem˜1, but where the size of the sketch depends on geometric properties of the manifold, most importantly the intrinsic dimension, and not on the collection of points being sketched. This means that for low-dimensional manifolds of a potentially high-dimensional ambient space, one can construct a small JL transform for an arbitrarily large number of points. The size $k$ depends only logarithmically on the ambient dimension $N$, and in fact, this dependence can be removed completely as shown in subsequent works [17]. This ability to construct JL transformations for manifolds is key to our main result, which we present next.
@@
Define $\mathcal{L}$ as the squared $\ell_{2}$ (i.e., unnormalized MSE) loss. For a fixed mesh $\bm{X}$ and time $t\in[T]$, with parameters $\bm{\theta}\in\mathbb{R}^{N}$, suppose the map $\bm{\Phi}:\mathbb{R}^{N}\to\mathcal{M}\subset\mathbb{R}^{n\times c}$, defined via the neural network as $\bm{\Phi}(\bm{\theta}):=\bm{\Phi}(\bm{X},t;\bm{\theta})$, maps to the Riemannian manifold $\mathcal{M}$. Further, suppose the sketch $\bm{S}_{t}:\mathbb{R}^{n\times c}\to\mathbb{R}^{k\times c}$ is a $(\epsilon,\delta)$-JL transform on the shifted manifold $\mathcal{M}-\{\bm{U}_{t}\}$, then, $\forall\bm{\theta}\in\mathbb{R}^{N}$, the following holds with probability at least $1-\delta$:
@@
The proof of the second inequality is motivated by the informal proof of Theorem 2.14 in [55]. Also, note that we construct a JL transform for the shifted manifold $\mathcal{M}-\{\bm{U}_{t}\}$, which is just a convenience to simplify the proof, but does not introduce any complexity to the sketch itself. The shifted manifold will have the same geometric properties as $\mathcal{M}$ that impact the construction of the sketch.
@@ 3 Experiments
We conduct our experiments on the datasets listed in Table˜1, which highlight several potentially challenging or important attributes. The Ignition dataset [^6] An internal dataset that was also used in [12]. describes a fully resolved gas wave front on a 2D uniform grid that transitions from a transport phase to a steady state jet. The Neuron dataset describes a 3D diffusion process in an unstructured, non-Cartesian branching neuron tree [32]. The Channel dataset, from the Johns Hopkins Turbulence Database [21, 33], describes 3D turbulent flow on a non-uniform grid. It has been trimmed in space to a $64^{3}$ volume and in time to the first 500 snapshots. We define the notion of a “sample factor” as the relative size of the sketch to the mesh size for a given dataset, i.e., $100\cdot k/n$.
@@
Equation˜6 provides the snapshot relative loss function we use for training, Eq.˜7 defines the full dataset relative Frobenius error, and Eq.˜8 defines the average snapshot peak signal to noise ratio (PSNR). The latter two of these three serve as our primary performance metrics.
@@
${\color[rgb]{0,0,0}\text{PSNR}(\bm{U},\bm{\widetilde{U}})=\frac{1}{T}\sum_{t=1}^{T}\frac{1}{C}\sum_{c=1}^{C}20\cdot\log_{10}\left(\frac{\max_{i}(\bm{\widetilde{U}}_{t})_{i,c}}{\|(\bm{U}_{t})_{:,c}-(\bm{\widetilde{U}}_{t})_{:,c}\|_{2}}\right).}$ (8)
@@
Our work is implemented in PyTorch [2] and available as an open-source GitHub repository [49]. All training is conducted on a single GPU (A100 or L40S) using the RAdam optimizer, a learning rate of $10^{-4}$, and in single precision (i.e., 32-bit floating point) [26] for both the parameters and data. Given that our goal is to compress the given data, we do not seek to generalize to unseen data; thus, overfitting is not an issue. Our method is fully in situ, so we perform no pre-processing of the data. For all in situ results, the size of the full buffer is always $T_{f}=1$ (only one full snapshot available at a time), and the size of the sketch buffer is $T_{s}=T-1$ (one less than the full time horizon). We do not require sketching the final snapshot because the training loop ends after the last full snapshot has been encoded. The batch size of the sketch buffer and the number of epochs per full snapshot vary between datasets, but are in the range 25–45 and 300–500, respectively. The sketch regularization weight from Eq.˜4 was manually tuned to $\lambda=5$ and fixed for all experiments.
@@ 3.1 Model Architecture
As alluded to previously, we elect to use a hypernetwork approach for constructing our neural compressor. Practically, this results in two networks working together to map space-time coordinates to a reconstruction of the PDE solution field. Figure˜4 displays the generic INR architecture we employ for both the hypernetwork and the target network. The “Hypernet” maps from the time coordinate $t$ to parameters and the “Target INR” ingests these parameters to define its layers which map from the space-time coordinates $(\bm{x},t)$ to the reconstruction $\widetilde{\bm{u}}(\bm{x},t)$.
@@
Within the sine-layer blocks, we use sinusoidal activation functions with the initialization developed in [50]. We also employ skip connections similar to those of [39, 46]. Thus, our networks are defined by two hyperparameters: the depth, i.e., the number of blocks, and the width of the internal linear layers. Because we are targeting the in situ setting, we cannot effectively normalize the data to ensure it is in some specific range, so no activation is used at the final linear layer. This ensures that the target network can produce any signal value. The lack of activation is also useful for the hypernetwork, as we construct the initial output of the hypernetwork to match the initialization of the target INR. To accomplish this, we use the strategy of [24], wherein the bias of the final linear layer in the hypernetwork is set to the exact values given by the initialization of the target INR, and the weights of the last linear layer are scaled by a relatively small factor.
@@ 3.2 Results
Table 2: Performance metrics on test datasets Ignition, Neuron, and Channel at 5%, 1%, and 2% sketch sample factors (where relevant) — which equates to equivalent offline storage costs for 22.5, 5, and 10 full snapshots — respectively. Metric statistics are reported as mean plus-or-minus standard deviation from five identical trials. The numbers in the parentheses of the first row of each table are compression ratios.
@@
The results in Table˜2 also show that our in situ methods, which combine training of the full snapshot with regularization from previous sketched snapshots, outperform the Offline-FJLT strategy (which just uses the sketched snapshots). The Offline-FJLT is conceptually simple to implement, since it does the sketching in situ and then trains offline. However, since it cannot train on the unsketched data, it incurs the $\frac{1+\epsilon}{1-\epsilon}$ error. Our in situ methods use sketched data but do not seem to incur this extra error, since the sketching is only intended to prevent forgetting, but the network can still train on unsketched data. This is also unsurprising, since for human memory, the task of learning something new is usually more difficult than the task of not forgetting something already learned.
Reported separately in Table˜3, we have also compared to another in situ baseline in the form of ZFP [35]. Here, each snapshot and channel is compressed separately through the zfpy software [10]. Comparable compression rates to what we obtain with INC are not achievable, but since ZFP is an error-bounded scheme, we can ask for similar RFE errors. At this similar quantitative performance, in terms of our RFE and PSNR metrics, ZFP yields two orders of magnitude lower compression rates.
Dataset Method Compression Rate RFE (%) $\downarrow$ PSNR (dB) $\uparrow$
[…]
Table 3: Baseline in situ performance of ZFP on test datasets. InSitu-FJLT results have been reproduced from Table˜2 for comparison.
In Table˜4, we explore how well the theory from Section˜2.2 aligns with our empirical results. Although unreported in Table˜2, we use Offline-FJLT and Offline-Baseline results for Eq.˜6 to compute the $\frac{1+\epsilon}{1-\epsilon}$ ratio from Theorem˜3, which we subsequently use to solve for $\epsilon$. We also compute principal component analysis (lPCA [19]) based estimates of the local manifold dimension for the INC networks used with each of the target datasets. This local dimensionality estimate yields a rough idea of the overall manifold dimension. We perform this analysis using the scikit-dimension package [4] by taking perturbations of size $10^{-5}$ around a nominal point of network parameters. This nominal state is obtained within the in situ setting after training on the first snapshot, i.e., the first time sketching is used.
@@
Figure 6: Reconstruction comparison of the Channel dataset on all three channel flow velocities $(x,y,z)$ components in top to bottom order) at snapshot $t=250$ for a constant $z$ coordinate in the center of the volume.
Per-snapshot relative error (Eq.˜6) curves for training and testing on the Ignition dataset are given in Fig.˜7. The training error is split according to Eq.˜4 into the two components $\mathcal{L}_{\text{full}}$ and $\mathcal{L}_{\text{sketch}}$ corresponding to the two buffers. Thus, the “Train Error (Full)” can be viewed as the instantaneous error at the current snapshot, while the “Train Error (Sketch)” is the averaged error over previous sketched snapshots. The “Test Error (Full)” is simply the instantaneous error for a fixed snapshot after the network has been trained. For early snapshots, before approximately $t=50$, one can observe that the testing error is significantly better than the training loss, indicating the sketched snapshots have also actively been used for learning instead of only preventing forgetting. While this is not necessarily problematic, it may suggest we could alter our training strategy, e.g., by using smaller sketches, to yield a more optimal process.
Figure 7: Training and testing error Eq.˜6 on the Ignition data for the “InSitu-FJLT” case. Training error is taken at the end of the optimizer steps for a given snapshot and split according to Eq.˜4. Data has been smoothed via convolution with a Gaussian.
@@
One may ask whether the inclusion of the hypernetwork in our overall INC architecture, shown in Fig.˜4, is necessary? Indeed, sketching as a regularization is just as applicable to a time-dependent INR, which we have found to be just as expressive in the offline setting (see Fig.˜2). To probe this question, Table˜5 presents comparison results on the Ignition dataset in the in situ regime with and without a hypernetwork. In the latter case, the hypernetwork branch seen in Fig.˜4(b) is simply eliminated. Removing the hypernetwork results in fewer total parameters, so the “INR” models are appropriately upsized via an additional SIREN block and larger linear layer widths to ensure approximately the same compression rate. We include results for both our subsample and FJLT sketching options. We observe that the addition of a hypernetwork significantly increases performance regardless of the sketch.